OpenStudy (anonymous):

x^2 - 1/36 Factor this, is the answer: (x+1/6)(x-1/6) thanks

OpenStudy (anonymous):

You can view this also as a difference of 2 squares.

OpenStudy (anonymous):

\[(\frac{1}{6})^2=\frac{1}{36}\]

OpenStudy (anonymous):

Your answer is right

OpenStudy (anonymous):

ok thank you :)